OpenStudy (anonymous):

Indicate the equation of the given line in standard form. The line that contains the point Q( 1, -2) and is parallel to the line whose equation is y - 4 = 2/3 (x - 3)

OpenStudy (anonymous):

y - 4 = 2/3(x - 3) y - 4 = 2/3x - 2 y = 2/3x - 2 + 4 y = 2/3x + 2 As you can see, the slope of this line is 2/3. And since you want to find a parallel line, then the slope will be the same....2/3. y = mx + b.....using slope 2/3 and points (1,-2) -2 = 2/3(1) + b -2 = 2/3 + b -2 - 2/3 = b -6/3 - 2/3 = b - 8/3 = b y = 2/3x - 8/3 Lets remove the fractions by multiplying by 3 3y = 2x - 8 Now we will put this in standard form -2x + 3y = -8 (multiply by -1 to make x positive 2x - 3y = 8 <-- this is your parallel equation

OpenStudy (anonymous):

Thanks